I need your help in finding the solutions to x^2-sqrt(5)x+ 5/4=0.
What I have so far inside the square root is -5-(4)(1)(5/4).

Answer by fractalier(6550)

For x^2 - sqrt(5)x + 5/4 = 0,
we have
x = (sqrt(5) ± sqrt(5 - 5)) / 2
x = sqrt(5) / 2
a double root...